Updated Jul 26, 2022 - 22:00
The following radiation cooling model well explained the cooling curve of the test mass (until ~150K)
where dQ/dt is the heat removed from the test mass, A is the surface area of the test mass, is the Stefan-Boltzmann constant, T_SH and T_TM are the temperatures of the surrounding shield and the test mass.
Can we extract any information from this "0.15"?
I borrowed "Cryogenic Heat Transfer (2nd Ed)" by Randall F. Barron and Gregory F. Nellis (2016) from the library.
P.442 Section 8.5 Radiant Exchange between Two Gray Surfaces can be expressed by Eq 8.44
where T_i is the temperature of objects 1 and 2. For us, OBJ1 is the test mass and OBJ2 is the shield. A1 is the surface area of A1. F_1,2 is the view factor and is unity if all the heat from the OBJ1 hits OBJ2. (It is the case for us.)
is an emissivity factor.
The book explains some simple cases in P 443:
Case (a): If OBJ2 is much larger than OBJ1, where the e_i is the emissivity of OBJi. This means that the radiated heat from OBJ1 is absorbed or reflected by OBJ2. But this reflected heat does not come back to OBJ1. Therefore the radiative heat transfer does not depend on the emissivity of OBJ2.
Case (b): If OBJ1 and OBJ2 has the same area, . The situation is symmetric and the emissivity factor is influenced by the worse emissivity between e1 and e2. (Understandable)
Case (c): For general surface are ratio, . OBJ2 receives the heat from OBJ1 and reradiates it. But only a part of the heat comes back to OBJ1. So the effect of e2 is diluted.
For our case, OBJ1 is the Si mass with DxH = 4in x 4in, while the shield is DxH = 444mm x 192mm. A1/A2 = 0.12.
We can solve this formula to be Fe=0.15. e1 = (0.147 e1)/(e2-0.0178).
Our inner shield has a matte aluminum surface and is expected to have an emissivity of ~0.07. This yields the emissivity of the Si test mass to be e1~0.2
How about the sensitivity of e1 on e2? d(e1)/ d(e2) = -0.95 (@e2=0.07).
Can Aquadag increase the radiative heat transfer?
Depending on the source, the emissivity of Aquadag varies from 0.5 to 1.
e.g. https://www.infrared-thermography.com/material-1.htm / https://www.mdpi.com/1996-1944/12/5/696/htm
It seems that painting Aquadag to the test mass is a fast, cheap, and good try.
[Returned] Brought one HAM-A coil driver (D1100687 / S2100619) and one Satellite Amplifier (D1002818 / S2100741) from the 40m
Also brought some power cables.
Brought ~1m of 0.0017" (~43um) misical wire. This will make the tension stress be 341MPa. The safety factor will be ~7.
- Temperature Log updated 2021/7/23 12:00 Heating Ended
- Assuming reaching the room temp at ~30hrs and heating power saturated at 22W: Predicted heat injection 30*3600*22 = ~2.4MJ
Update from Stephen
- Note that we can check logging accuracy against the snapshot (timestamp 20210723_1113).
If my math is correct, this would be time = 37.35 38.35 hours
Update from KA
=> The corresponding time in sec is 138060 sec
The raw data line for the corresponding time is:
138016.839614, 295.805, 306.678, 302.518, 312.401, 0.000, 0.000, -0.001, 0.621, 0.622, 1.429, 0, 0, NaN, NaN, NaN
The values on the photo 295.806, 306.677, 302.518, 312.401 ==> Well matched. Victory!
[Stephen and Koji for discussion / Koji for the execution]
1. Temperature Trend
See [QIL ELOG 2611] for the updated temp log and the cooling model.
Considerations for the next cycle:
-> How can we accelerate the cooling? It seems that the table cooling is conduction limited. Improve the cold head connection.
-> We want to move the RDTs
-> How can we improve radiative cooling?
2. Oplev Trend (Attachment 1)
Sum: The beam has been always on the QPD (good). See also Attachment 2
X&Y: In the first few hours the beam drifted in -X and then +X while Y had slow continuous drift in +Y. ~11hours later sudden drift in -Y and totally saturated. Also -X saturation observed @~16hrs. Again +Y drift was seen @~25hrs. The totally saturated in -X and +Y.
They may be related to the drift of various components with various cooling time scale.
Visual check: ~2mm shift in X&Y is visually observed. Attachment 2
-> How can we quantify the drift? What information do we want to extract?
3. OSEM and the magnet
The magnet is intact. And the suspension seemed still free after cooling (Attachment 3)
Significant misalignment was not visible. No visible damage by cooling was found. The coil is alive and the PD/LED are also intact. Fluke showed that they are still diodes, but their function was not checked.
The coil resistance changed from 16Ohm -> 4.2Ohm. For the 16Ohm, 2 Ohm was from the wire. Let's assume we still have 2Ohm overhead -> The coil R changed from 14->2.2. This corresponds to the coil temperature of the order of ~100K. This is not so crazy.
Some actuation current was applied to the magnet. For this test, the oplev was realigned.
First, some ~300mA current pulses were applied to the coil. The ringdown of the yaw mode was visible. Then the DC current of 100mA was applied. This didn't make visible change on the spot position but the data showed that there was a DC shift.
-> We prefer to have a softer suspension for the next test.
4. CTC100 logging
During the cooling we kept having inaccurate data logged compared with the displayed data on the screen of CTC100.
As soon as the cooling logging was stopped, telneting to CTC100 was available. So, I telnetted to the device and sent the data transfer command ("getOutput"). Surprisingly, the returned values agreed with the displayed values.
So my hypothesis is that somehow the data strings are buffered somewhere and gradually the returned values get delayed. From the behavior of the device, I imagined that the fresh telnet connection gives us the latest data and there is no buffering issue.
So I tweaked the data logging code to establish the telnet connection every time the values are asked. The connection is closed after the every data acquisition. I like this as we can also make the test connection between each data acquisition points, although I have not tried it yet. The code is in the same folder named ctc100_controller_v2.py
Now I thought that I did all I wanted to do this evening, so the heater was turned on at ~20:50, Jul 21. The heating power saturated at 22W, which is the set limit.
Yesterday I came in the QIL and performed an express kidnapping of the 2um in fiber AOM (Brimrose) and the 5 W RF amplifier that was hooked to the RF in port (though it seems it saturates at ~ 600 mW from past elogs). I will test it with the 1419 nm ECDL fiber pickoff port to see that it works and if it doesn't I will reinstall it in the 2um testing facility.
Updated the model the latest log data with cooling prediction
Radiative cooling already gives us a good agreement with the measured temp evolution for the test mass. The conductive cooling is not significant and does not change the prediction.
Updated the plot with the new data (2021/7/21 12:30PM)
A naive cooling model was applied to the cooling curve.
A wild guess:
- The table temp is the same as the test piece temp as measured on 2021/7/9
- The inner shield temp is well represented by the table temp
- The specific heat of Si is almost constant (0.71 [J/(g K)] between 300K~200K
The curve was hand-fitted by changing the emissivity of the inner shield and the silicon mass. I ended up having the same values for these to be 0.15.
Surprisingly well fitted!
The conductive cooling through the wire does not fit the cooling curve, although the quantitative evaluation of the wire conductivity needs to be checked carefully.
Stephen shared attachments 2 and 3, which contain insights on the wire used to hang the Si mass. .017" diameter Music Wire from California Fine Wire, 2004 vintage, borrowed from Downs High Bay.
Temp Log on Jul 19 2021 17:20
I wonder what is the heat transfer mode for the test mass right now. Radiative? or Conductive through the wires?
Uh oh, review of the cooldown plot from the previous cooldown (QIL/2603) shows workpiece temperature of ~92 K at conclusion, while a temperature of 65K was observed in the CTC100 readout (Attachment). The logging of the warmup is consistent with the CTC100 image, as the logging started a few minutes after the warmup was started, and the warmup "5 minutes after starting" temperature of ~ 71 K is a practical temperature.
Seems to be something weird going on here, we will need to have Radhika take a look on her return (and continue taking photos of the CTC100 whenever we stop by).
Temperature log for the first 2 hours (Attachment 1)
I wonder why the temperatures displayed on CTC100 and the ones logged are different...?
We started cooling down of the test mass.
- Stephen vented the chamber at 2PM. An optical port was moved to see the OSEM from the back.
- Brought DSub crimp sockets from the 40m. We picked up 3x 1m LakeShore WCT-RB-34-50 (twisted silver-plated copper, 34 AWG with Teflon insulation). The ends of the wires were dangled so that crimping is possible. A single wire resistance was measured to be ~1Ohm at room temp. (Attachment 1)
- OSEM pin out / backside view (cable going down) (Attachment 2)
| o o o |
| o o o | Wire
^ ^ ^ ^ ^ ^---PD K ---- R3
| | | | |-----PD A ---- B3
| | | |-------LED A ---- B2
| | |---------LED K ---- R2
| |-----------Coil End ---- B1
|-------------Coil Start ---- R1
Twisted Pair 1: (R1&B1) with 1 knot at the feedthru side
Twisted Pair 2: (R2&B2) with 1 knot at the feedthru side
Twisted Pair 3: (R3&B3) with 1 knot at the feedthru side
| o o o |
| o o o | Wire
^ ^ ^ ^ ^ ^---PD K ---- R3
| | | | |-----PD A ---- B3
| | | |-------LED A ---- B2
| | |---------LED K ---- R2
| |-----------Coil End ---- B1
|-------------Coil Start ---- R1
Dsub feedthru in-air pinout (Mating side)
1 2 3 4 5
\ o o o o o /
\ o o o o /
6 7 8 9
1 2 3 4 5
\ o o o o o /
\ o o o o /
6 7 8 9
Pin1 - Coil Start
Pin6 - Coil End
Pin2 - LED K
Pin7 - LED A
Pin3 - PD A
Pin8 - PD K
Pin2-7 Diode V (with Fluke) 1.18V (Pin2 black probe / Pin7 red probe)
Pin3-8 Diode V (with Fluke) 0.7V (Pin3 red probe / Pin8 black probe)
- OSEM pin out / backside view (cable going down)
Suspension installation (Attachment 3)
- The sus frame was moved into the chamber
- We measured the test mass dimension before installation: L 3.977" D 4.054"
- The attached mirror size is 1"x1" made of SUS #8 (?)
- The mass was suspended. The height / rotation of the mass was adjusted so that the reflecting mirror is visible from the oplev window and also the OSEM magnet is visible from the OSEM window.
- The OSEM was placed on an improvised holder. (Attachment 4)
- ...Just the usual oplev installation. Adjusted the alignment and the return beam hits right next to the laser aperture. This beam was picked off by a mirror and steered into a QPD. (Attachments 5/6)
- The lever arm length is ~38" (960mm) -- 9" internal / 29" external
- The oplev signal is shaking so much and occupying ~50% of the full scale. Added a lens with f=250 to make the beam bigger, but the improvement was limited.
- Started ~8:30PM?
- Wired 3 BNC cables from the table to the DAQ rack. CHX/Y/S are connected to ADC16/1718ch.
- The real-time processes seemed dead. Looked at [QIL ELOG 2546] to bring them up. TIM/DAQ error remains, but the data stream seems alive now. Leave it as it is.
- Temp Logging started. Filename: temp_log_cool_down_20210716_2255.txt
- Cryocooler turned on. ~10:55PM
- Confirmed the cold head temp was going down. The cold head temp is 75K at 0:30AM
- An example photo was taken from the rear window. The attempt with 40m's Canon failed. Attachment 7 was taken with KA's personal compact camera with a smartphone LED torch. The gap between magnet and OSEM is highly dependent on the view axis. So this is just a reference for now.
[Stephen / Koji]
Bonding work for the prep of the preliminary suspension test
- 1" sq mirror-ish polished SUS piece was bonded to a face of the silicon mass. We chose the location right next to a line on the barrel. (Attachment 1)
- The mass was flipped with two more same thickness pieces used for the spacers to keep the mass horizontal.
- A pair of an OSEM and dumbbell-magnet was brought from the 40m (courtesy by Yehonathan). The magnet was glued on the mass at the opposite position of the attached mirror because the optical ports are going to be arranged to share an axis. A piece of cryo varnish was also painted with a piece of cigarette paper at the center of the mass so that we can attach an RTD. (Attachment 2)
Next Things To Do (Attachment 3)
[Stephen Koji Radhika]
Stephen and Radhika worked on the cooling down and warming up of the cryostat with the cold head RTD attached using a spring-loaded screw. No other configuration changes compared to QIL/2599. Here are the temperature log plots. Photos of spring clamped RTD are outstanding, but the clamp is the same as the workpiece pictured in QIL/2599/Attachment 12.
Brief summaries of the last week's progress and the coming week's plans (plots will be posted soon!):
- progress Friday 09 July: Opened the cryostat up at the cold head, and attached an RTD to the cold head with a spring clamp (instead of relying purely on the cryo varnish).
- progress Monday 12 July: Found 65 K workpiece temp and 63 K cold head temp. RTD was apparently held successfully by spring clamp, and we will continue to collect cold head temperature in future runs. Warmup was started, with old data collection completed (cooldown_20210709) and new data collection commenced (warmup_2021_07_12). Note that warmup started at 1:14 pm, and it took me ~ 5 minutes to stop and restart the script to changeover to the warmup data collection.
- table plan Wednesday 15 July: Complete in-air optical layout. Make one flat face of Si mass reflective.
- chamber plan Thursday 16 July: Open up main volume and drop in frame with Si mass. Connect RTDs. Start cooldown. Confirm cooldown is going ok (optical alignment, especially), and revert if necessary before things get too cold.
- table plan Friday 16 July: Maybe measure stuff, maybe better to wait till coming week and use controlled heating to hit different temperature setpoints.
*Takeaway*: The current 1D cooling model is getting closer to matching our observed cooling trends, mainly in the lower temperature limit. The predicted time constant is still much smaller than we are seeing in reality (by about a factor of 3), but this can potentially be improved by revising specific heat values and/or dimensional estimates for chamber components.
The model uses the known cooling power of the cold head [attachment 2] and considers radiative heat from the outer shield, baseplate (bottom lid), and mylar wrapping around braid. I increased the complexity of the script by solving a system of ODEs (for braid and coldplate temperature) simultaneously instead of assuming the temperatures are equal at all times, and solving only 1 ODE. This resulted in the model's lower temperature limit prediction matching our observed data, at ~66 K.
The model still predicts a much smaller time constant than we are seeing. This is affected by specific heat values for Cu and Al, along with dimensional estimates of the coldplate and braid (AKA how much mass is being cooled). It is possible that these values are being underestimated in the model, which would lead to the smaller time constant. Currently the model uses constant values for the specific heat of Cu and Al (room temperature). But since specific heat increases with temperature, accounting for temperature dependence would lower the specific heat values and shift the model in the opposite direction (towards an even smaller time constant). Therefore I suspect the model is underestimating the mass of the coldplate, though I am unsure if this would completely correct the discrepancy.
If the term (specific_heat * density * volume) of the coldplate (Al) is increased by a factor of 4, the model resembles the data well [attachment 3].
I wrote a python script to extract temperature data from the CTC100 via ethernet, for monitoring cooldown/warmup of Megastat. This is intended to replace USB data extraction, which requires the user to manually insert/remove the stick and plug into a computer.
The script queries the CTC100 every ~60 seconds for the latest temperature values (the frequency can be supplied as a parameter, but default is 60s). The script writes line-by-line to a .txt file and also plot the outputted data once collection is terminated.
Here is a gitlab link to the script: https://git.ligo.org/voyager/mariner40/-/blob/master/CryoEngineering/ctc100_controller.py. It is also found on the QIL workstation at /home/controls/CTC100/ctc100_controller.py. To run from the workstation, open terminal to /home/controls (home). Then:
python ctc100_controller.py --filename='tutorial'
Here, 'tutorial' stands in for the desired filename for the outputted data. The script will start pulling data and will print each line to the terminal. It will continue printing and logging the temperature values until the user hits Ctrl+C in the terminal. This will terminate the script and output the final data file. The file is saved as a .txt file in /home/controls/CTC100/data.
I've attached cooldown data from 7/1-7/4, which was extracted via ethernet from the CTC100. Initially I was having issues getting the curves smooth; the data seemed very choppy and I assumed it was a device or ethernet issue. Turns out the csv file outputted by my script would cast all floats to integers if I opened it and saved, which was very strange. I'm still not sure why this happens, but my workaround is to not edit the csv at all after it is outputted.
The workpiece temperature stabilized at ~65.8 K, just below 66 K (last cooldown). This makes sense given the few minor tweaks we made in between (removed shorting Al foil, adjusted mylar / peek sheet insulation). The time constant is also consistent (~36.4 hrs).
Unfortunately the RTD on the cold head appears to have detached yet again. It seems to have happened around the same time as the last cooldown. Stephen suggested replacing the current varnish bonding with a spring clamp, like what holds the workpiece RTD in place.
I switched off the cryo cooler at 12pm today (Tue), so the chamber should be ready for venting Thursday or Friday. The tentative agenda is to add the spring clamp to secure the cold head RTD and cool down again (we will discuss any additional tweaks). I think I have an idea of why the current model predicts a much smaller time constant than we are seeing - I plan to work on this more this week. Hopefully once the cold head RTD stays fixed for an entire cooldown, this data can help tweak the model as well.
On Tuesday 6/29, we opened up Megastat after a weekend of warmup. Our goals were to 1) check on the RTD attached(?) to the cold plate to confirm that it had de-bonded, and 2) to check on the foil and mylar inside the chamber and around the copper braid.
Indeed the tip of the cold head RTD had detached [attachment 1], but the mylar "cap" I created for insulation held up well. Inside the chamber, we noticed some shorting between aluminum foil on the inside of the outer lid and the outside of the inner lid [attachment 2]. Also, there was a gap in mylar around the copper braid, so some of the braid was visible. We also noticed that the strip of peek sheet inside the opening of the inner shield had slipped out of place, exposing the copper braid to the aluminum [attachment 8].
We rebonded the loose RTD to the cold head [attachment 3] and re-inserted the mylar "cap" [attachment 4 and 10]. Stephen suggested we add an additional piece of mylar to cover the portion of the cold head directly above the braid [attachment 5].
To prevent shorting inside the chamber, we trimmed the excess aluminum foil wrapped around the non-coated sides of the inner and outer shields [attachment 6]. We adjusted the mylar sheets around the copper braid to cover the previous gap, and used a mirror to ensure the bottom of the braid was covered [attachment 7]. We also realigned the peek sheet strip with the opening of the inner shield [attachment 9].
We reattached the heater to the workpiece [attachment 11] and the workpiece RTD with the spring clamp [attachment 12]. We placed back the inner and outer shield lids [attachment 13-14], closed up the chamber, and pumped down. Since I have been working on remotely extracting temperature data from the CTC-100, we chose to wait until the script was ready before turning on the cryocooler.
On Thursday 7/1, I got the temperature extraction script ready and turned on the cryocooler at ~5:15 PM. Fun addition: attachment 15 shows successful communication to the CTC-100 (I sent a command to display "hello" on the screen :D)
On Friday 7/2, I plotted the temperature trends so far [attachment 16]. As of 12pm Friday, the workpiece is at 177K. We will let the chamber cool over the weekend and see where we are on Tuesday.
Radhika and I started discussing in detail the items needed to transfer the 2um PD setup into the Cryo Vacuum Chamber.
To interface with the PD setup, it seems that we will need want to use the feedthrough already in use on the IRLabs cryostat. This square-flanged feedthrough appears to host the in-vacuum cabling. The in-air cabling would ideally also be reused.
To implement this, we will need to modify a blank 2.75" conflat flange (in hand in QIL boxes on wire shelf) with a the following features:
I will take care of this part through the PMA shop.
The Cryo Vacuum Chamber windows are Thorlabs VPCHW42‐C (ref. D2000310-v1). The broadband AR coating is rated for 1050-1700 nm (here's the data sheet), with a steep climb in AR coating reflectance. There is also fused silica substrate transmission to worry about. Perhaps the easiest way to proceed would be to repeat Koji's Transmission test (ref. QIL/2458) for the windows already in use on the Cryo Vacuum Chamber, and see if we can accept ~10% loss (or more).
Else, perhaps we scavange the IRLabs cryostat and create another custom blank CF modification?
Or buy a new, suitable window with CF interface? (<---- preferred option)
Radhika and I were able to connect to the temperature controller using telnet. We simply reassigned the IP address of the temp controller to fit in the IP address range of the router/network. i.e. if the router is 188.8.131.52 then the temp controller might be 184.108.40.206. This shows that the router and network are working correctly again.
I was able to connect to the router (the one on top of the switch in the rack next to the orange tool drawers) and reconfigure it using the same internal IP as found labeled and mentioned in . The router now has a different external IP that I wrote on the router (twice). The network/router and switch is now working in the lab.
Update: It turns out that the router reset itself somehow. so any settings/ port forwarding that was set up before has probably been wiped out and would need to be set up again.
SSHing into the QIL needs to be set up again.
I've attached cooldown data from 6/21-6/24. Unexplicably, the CTC-100 stopped logging temperatures at ~50 hours after the start of cooling, or around 5pm Wednesday. The red logging indicator was definitely on at 3pm Wednesday, but it was off when I checked in Thursday afternoon. It must have somehow been disabled in between. On the bright side, the temperature of the workpiece stabilized at 66K. I had checked in at 53 hrs, 62 hrs, and 64 hrs and noted the temperatures by eye (green data points on attachment 2).
The time constant for the workpiece is ~ 36.4 hrs, about the same as the previous run (tau = -40/ln(100/300)). While no improvements were made on this front (cooldown still very slow), it stabilized 6K below the previous run.
The RTD at the cold head recorded a shoot-up in temperature right at the beginning of cooldown, which makes Stephen/me think that it debonded from the cold head. This temp reading would have been informative for our model, but we will verify it's status when we open up. We can find a better bonding mechanism for the next cooldown.
[Update] I let the chamber warm up over the weekend, and brought it up to room pressure on Monday 6/28. Next steps are to open up and investigate the cold head RTD and make improvements to bonding mechanism, if necessary. We will assess any other improvements (model driven) before the next cooldown.
I accidentally powered down and restarted the TP-link safe stream router and now nothing will connect to it. Not even things that are plugged into it. I can't even connect to the router's setup screen. I have no idea how to solve this issue short of resetting the router. I don't know how it was configured before so I will not reset it because I don't know how it was configured.
I brought DAC2ADC_test.py code from the 40m to test in the QIL. I added a parser arg for the matrix root name (see attached code). I am running into the same problem as I did at the 40m where the channels seem to be locked to their values. I attached (attachment 2) the results text file that shows that for all inputs the outputs are the same. In the GDS screen the ADC is in red but I'm not sure how it got there or how to fix it.
run the test with these commands in the QIL
$ cd IanMacMillan
$ python DAC2ADC_Test.py -e 'C4:TST' -m 'cdsMuxMatrix'
One improvement that we could make to the code is have it zero out all of the matrix elements before it starts.
Background: There are currently 4 available RTDs for the large cryo chamber (henceforth named Megastat). We originally had 1 for the workpiece, 1 for the heater, 1 for the inner shield, and 1 for the outer shield. During the last sessions with the chamber open, the RTDs were moved around. Now there is 1 for the workpiece, 1 for the baseplate (bottom lid), one for the outer shield, and 1 for the cold head. These locations are marked in the diagram (attachment 9).
On Wednesday 6/16, we opened up the chamber and removed the shields and coldplate. We added foil to the inner surface of the bottom lid (attachment 10) and attached an RTD there (we forgot to take a picture). Stephen made a new foil collar, and we decided to push it against the chamber walls as a better alternative to having the foil touch the coldplate (attachment 6). We added an extra layer of aluminized mylar to wrap the copper braid, and we fed another RTD through the braid tube to be attached to the coldhead. We lastly placed the coldplate and shields back in place.
On Friday 6/18, I attached the remaining RTDs (coldhead, outer shield, workpiece). We decided to remove the aluminum foil previously covering the coldhead, due to of fear of shorting to the tube wall. I taped several pieces of mylar together to cover the coldhead and insulate it from the wall (attachments 1-3). I placed the mylar contraption into the T (attachment 4) and then closed the bottom flange. I placed back the shield lids and the main lid of the chamber (attachments 7-8). I tried to pump down, but the pressure was stabilizing on the order of 1e-1 torr. Unsure of why the pressure wasn't decreasing, I turned off the pump and left for the day.
Today 6/21, I touched base with Stephen and we realized I forgot to replace the copper gasket in the bottom flange of the T. I then unscrewed and replaced the gasket. I pumped down and the pressure reached on the order of 1e-4 torr, so I proceeded with cooling. After a few hours I could see that the cold head RTD was reading a temperature around 180K. We will extract and plot cooling data late tomorrow or Wednesday.
Lastly, a chamber diagram is attached (attachment 9). The 4 RTD locations are marked by a red 'x'. The chamber components are numbered in blue (detailed below):
3. Inner shield
4. Outer shield
5. Copper braid (wrapped with mylar)
6. Cold head
8. Baseplate (bottom lid)
RTD thoughts - we have just been using the sensors that were provided, without noticing their constraints or deficiencies.
Planning for next steps:
I was searching an I2 (Iodine) cells back to the days of the laser gyro.
I found a likely box at a very tricky location. Took the photos and returned to this tricky place.
2021/Jul The box was moved to the OMC lab (KA)
We hit another dead-end with leak hunting the IR labs dewer (we replaced screws and helicoil on the valve connection but there is still a big leak). We cleaned the flange and O-ring with isopropanal and replaced the threads with helicoil but still get the same sort of leak where we only hit 1E-2 Torr after 5 minutes of pumping and stablize around 1E-3.
After turning off the pumping station, the pressure rose quickly to 1Torr (in roughly 10 minutes or so).
[Stephen, Aidan, Wednesday 09 June]
Summary and Plan:
Troubleshooting steps taken:
Summarizing heater trends from Wednesday, 6/9. Reiterating from post , I ran the CTC-100 temp controller's auto tune routine to adjust PID coefficients for the heater. I then set the setpoint to 123K (operating temperature); the controller takes in the workpiece temperature as feedback. The heating data for this period is attached.
It took under 30 minutes for the temperature to rise from 72K (current cooling limit) to within a degree of the setpoint, 123K. The power delivered to the heater (bottom plot of first attachment) stayed below 25W, the limit we hardcoded. The workpiece temperature rises smoothly and plateaus around the setpoint, without significant overshooting. The controller holds the setpoint temperature pretty well thereafter. The second attachment is zoomed in on the workpiece temperature alone.
This run served as a test of the temperature controller's stability at the desired setpoint. Moving forward, we will continue to improve the cooling capacity of the chamber guided by our model. Once we optimize how cold we can get, we now know the heater can hold us at the desired temperature setpoint.
Code for simple heat transfer modeling can be found here: https://git.ligo.org/voyager/mariner40/-/blob/master/CryoEngineering/qil_simple_heat_transfer.ipynb
My original 1D cooldown script modeled conductive cooling along the copper braid as: Pcool = k * A/L * (T - T_set); where Pcool is the cooling power, k is the thermal conductivity of copper, A is the cross-sectional area of the braid, and L is the length of the braid. I changed the code to instead use the tabulated power vs. temperature points for the CH-104 coldhead, taken from Paco's script qil_heat_estimate.ipynb. The first figure compares the interpolated curves from the tabulated values (at 50Hz and 60Hz operation), to the original conductive transfer model (50K setpoint). The original conductive power-temp relationship is linear, which overestimates the cooling power at high temperatures. Switching to the tabulated points results in more realistic model. Moving forward, I intend to use the 50Hz interpolated curve.
The script considers radiative heating to the coldplate from the the chamber bottom (rough aluminum) and the outer shield (coated in aluminum foil). It assumes over a long period of time that the inner shield and coldplate temperatures are equal.
The second figure shows the results of this model alongside the actual coolddown data extracted from the CTC-100. It is clear that the model is not accounting for additional radiative heat sources that would explain the slower cooldown and higher final temperature. Adding in model complexity is my current focus.
Today I attempted to auto tune PID coefficients for the heater, so that we can reach and maintain a setpoint of 123K with appropriate ramp-up. The workpiece was around 72K originally. For auto tuning, I set the max power for the heater to 25W. I adjusted the lag time to 30s, and changed the setpoint to 72K so that the tuning response measured how stable the system is when being perturbed from our set point. The auto tune process ran without an error; however, by default the tuning mode switched to "step" tuning, and I am not sure why this occurred. The tuning took roughly 5 minutes to complete. The final message is attached; the adjusted parameters were:
Lag: 125.8 s
Time constant: 271.9 s
I was expecting the adjusted output to be new PID coefficients, but I noticed that the PID coefficients did seem to change after this process. To begin warmup to 123K, I changed the setpoint to 123K and let the heater do its thing (the feedback temperature is set to that of the workpiece). The heater power stabilized to around 17W, and the temperature of the workpiece reached within a degree K of the setpoint within 30 minutes. I am letting the temperature hold at 123K overnight and plan to return tomorrow morning to check on it and extract the heating data for plotting.
Quick log establishing the maximum power for our thermal actuation:
Heater: HSA25100RJ from TE, unknown sourcing. Acetone wiping cleaned off p/n and markings from body, should engrave at next opportunity, but [Attachment 1] from many months ago shows the p/n. Note that this is not the current mounting configuration - [Attachment 2] is more similar to current mounting. Anyway, according to the datasheet (now added to the QIL wiki at Documentation > Manuals) this heater is rated for 25W and has a resistance of 100Ω.
Leads: unknown, and not super important unless we had tiny hair conductor - I am not in lab presently, but it appears from our connector (Lesker FTACIR19AC) that we must have 20-24 AWG,
Carrying the 7, the current through the Heater will be 0.25 A at max actuation, and the 20-24 AWG insulated copper leads will have plenty of ampacity for this load (plus, they are cooled, so normal current capacity considerations fly out the window a bit).
Conclusion: 25W actuation will be the limit that we will apply in the CTC100 temperature actuation routine.
From first trials yesterday, the response at ~20W (at starting temperatures around 80K) appears to be on the order of 1 degree per minute, which should be just fine for actuating to maintain a +/- 1 degree constant setpoint with static thermal loads. More to follow on trials implementing temperature control.
Note that the QIL Wiki points to the DCC (which contains a budget that was helpful resources to trace these purchases), the datasheets and other documentation, and also points to the QIL Cryo Vacuum Chamber photo album, which hosts the images below.
I extracted cooldown data from the CTC100 USB around 1pm today (~50 hours of cooldown). I've attached a log plot below. The heater RTD seemed to be behaving weirdly at the beginning, but soon stabilized and cooled as expected.
I estimated the time constant for the workpiece: 50 hr/ (300K - 90K) = 0.24 hr/K = ~860 s/K
I'll be curious to see the results of Radhika's thermal model - I am suspicious of this thermal strap contact to the base plate. It would be good if we could instead make a copper mating plate:
After Thursday's work, we resumed on Friday and lifted up the coldplate to access the collar and baseplate. Stephen added metallized mylar wraps to the peek cylindrical spacers [pic 9, 10]. He took out the cylindrical collar and baseplate (previously there to minimize contact between the colplate and chamber bottom) and replaced them with a foil collar [pic 13, note this is before pushing down the foil over the PEEK spacers]. We re-inserted the coldplate.
In order to wrap/insulate the copper braid, we inserted 2 sheets of metallized mylar into the vacuum tube to surround the braid [pic 11, 14]. We cut the mylar wrapping so that it did not short to the outer shield. We also confirmed that the copper braid is not shorting to the inner shield hole (there was clearance for the shield to be lifted before hitting the braid). The mylar extends to the coldhead, which we then wrapped with aluminum foil, making sure not to short to the walls of the tube [pic 12].
We switched the foil wrapping from the inside to the outside of the outer shield, so that any radiative transfer from the inner shield to the outer shield would be absorbed as much as possible (not reflected back). We placed both shields back and bolted down the copper braid loop and workpiece. We re-attached the RTDs, then placed in both shield lids (without bolting them down) and stopped for the day.
Today I checked on the chamber setup and closed up. I started the vacuum pump and it made abnormally loud noises, indicating something was off. The percentage sign continued to flash, indicating that the pump was not reaching 80% speed. I performed 2-3 power cycles, which did not solve the issue. We will pick up to debug the issue early this week.
[update 02 June 2021]
Pumpdown and cooldown were successfully started this morning - Radhika retightened the green leak valve and pumps started just fine. We will check trends on Friday and likely will allow cooldown to proceed over the weekend.
We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album.
This morning I vented and opened up the chamber to add aluminum foil and other insulation. Stephen's order of mylar sheets, peet sheets, and G10 rings came in today and I picked it up from Downs.
- cut narrow rings of 0.01'' thickness peet sheet and inserted into the main hole of the inner/outer shields, to prevent shorting to the copper braid [pic 6, 2, 3]
- added foil to the inside of the outershield and outside of the inner shield, poking holes for all viewports [pics 4,5]
⁃ unclamped the copper braid from the coldplate and the RTDs from all locations (shields, heater, workpiece mount) to prepare for lining baseplate with foil [pic 7]
Tomorrow we plan to wrap the shaft of the copper braid with metallized mylar and an additional layer of foil. I left both shields outside of the chamber so that tomorrow we are ready to remove the cold plate and add foil below.
Radhika, feel welcome to post full cooling data to this entry, or to your original - up to you!
I tried Krytox around the O-ring and also tightening the screws around the valve. The leaking persists at roughly the same rate.
I tried pumping down the JPL PD chamber to test the new PD at cryo temperatures. Unfortunately, the chamber can;t get past about 6E-3 Torr with the pump on. As soon as I turned off the pump the pressure rose to around 2 Torr over 20 minutes or so.
I extricated the chamber from the pedestals, flipped it and removed the bottom plate. I cleaned the O-ring with isopropanol and wiped down the mating surface on the chamber (also with iso). I replaced the plate and tightened the screws. Then I returned the chamber to the table and reconnected it to the vacuum system. I tried pumping down once again but I saw pretty much exactly the same situation as before (pressure bottoming out around 6E-3 Torr and then rising quickly again when the pump was turned off).
I guess it's possible that the O-ring is damaged - although I couldn't see anything obivous. We didn't mess around with the viewport (when we replaced the diode a few weeks ago) so I'm hoping there is no issue there.
Instant gratification McMaster sourcing (PO S519341, submitted earlier today)
Should be able to integrate some of these items soon - they arrive Thursday, and Radhika will check to see if anything works for the setup as-is, or make requests for Stephen to cut on Friday.
future Mylar source - https://www.professionalplastics.com/MYLARFILM
- update: Radhika called Professional Plastics and they said we cannot order metallized mylar online, but we can call them back and place the order. Stephen called later and Sarah said she had to get a quote from a supplier, and will be in contact. Thickness range: 0.001-0.014 in.
future PEEK source - https://www.professionalplastics.com/PEEK_SHEET-ROD-BAR: thickness range: 0.25-2 in. (units not specified, but I assume inches).
future PEEK source - https://www.boedeker.com/Product/PEEK-Virgin-Natural: thickness range: 0.062-4 in.
exp plot tip: If you use the "grid" feature of matplotlib and plt.semilogy(), the exponentials will look like straight lines, so we can just read off the time constants with a ruler.
Also, as we talked about earlier today, we should make some analytical estimates for the various heat loads, and also put them into the model.
For protecting from radiation, all of the surfaces which are NOT shiny-polished should get wrapped in something shiny (UHV Al foil, with the shiny side out).
I suggest wrapping with foil:
I attach here a photo of the radiative shielding of a Purple Pepper Plant (PPP), to reduce the radiative coupling to the environment. This prevents the soil from drying out in the sun so fast.
Today I pulled the cyro chamber cooling data from the temperature controller. Cooling started on Friday 5/21 around 1:45pm.
The final temperatures reached at ~3:15pm today were:
outer shield: 253 K
inner shield: 168 K
heater (off): 151 K
workpiece: 150 K
The temperature curves (see attached) seem to be leveling off, so I'm not sure things will get too much colder. In the meantime I've reinserted the USB to resume temperature logging.
[Radhika, Stephen remote]
After leaving the cryocooler's compressor running overnight, Radhika found all RTDs reporting room temperature. The noises coming from the compressor were normal, and all operating conditions were consistent with QIL/2504. However, the cryocooler was silent (valve motor not starting).
It turned out that the cryocooler power cable, unplugged during installation efforts (pictured below), had not been reattached. After turning off the compressor, plugging in the power cable, and turning on the compressor, the cryocooler began making normal noises and apparently operating normally. Radhika reformatted the USB drive collecting data from the temperature controller.
Cooldown began at 1:45pm today (Friday). We will check in again on Monday.
During these troubleshooting efforts, we referred to the Cryocooler and Compressor manuals, found at the QIL Cryo Vacuum Chamber Wiki.
I've returned the Keithley Source Meter unit
- The unit (Keithley 2450?2460?)
- A power cable
- A pair of banana clips
- the transistor test fixture & triax cable/connectors
Following up from Stephen's last post. Today I completed the outstanding tasks he outlined, with the exception of connecting a com cable to record and trend cooldown from the temperature controller. For today's cooldown we are still using the usb flash drive.
The RTD connected to the workpiece (spring clamp) took some wrestling to get stable temp readout, so I had to reclamp it. Other than that, I was able to close up the radiation shields' lids and the chamber lid straight away.
I initially tried pumping down but the backing pump was very loud, not reaching full speed, and pressure wasn't decreasing. Stephen realized the KF joint on the side of the chamber was never sealed up (just wrapped in foil) --> obvious major leak. We didn't have any blanks to seal it up, so I replaced the KF port with a blank flange at the conflat joint.
Ready to pump down once more, I ran into an error message from the turbo pump but performed a power cycle and it disappeared. Pumped down and reached a millitorr before turning on cryocooler.
I plan to pop by tomorrow morning before the cryo meeting and can share consequent updates then.
Good progress toward pumping down, with a setback (impact unknown while we reach out to Karthik).
The following is the list of remaining actions before we have cooldown data:
We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album. We will then collaborate to add some of the most interesting images to this log!
As of last Thursday (5/13), the new vacuum tubes had been selected and bolted in. We fixed the copper braid to the cryo-cooler: the braid was folded in half, with the loose ends bolted to the cryo-cooler and the folded end fed through the vacuum tubes into the chamber. The folded loop was bolted down to the baseplate. The copper braid was pulled tight and thus has no slack. Aluminum sheets were used to wrap frayed areas of the copper braid to prevent shorting to the tube walls, though this needs to be revisited. We also still need to address shorting to the inner/outer radiation shields.